OpenStudy (anonymous):

Can someone create for me an exponential function that increases gradually? :)

OpenStudy (amistre64):

take the inverse of a log the increases rapidly ....

OpenStudy (zehanz):

Isn't every exponential function increasing gradually? What is your definition of "gradually"?

OpenStudy (dumbcow):

exponential function \[f(x) = b^x\] for b>1 , it increases larger b will cause function to increase faster

OpenStudy (zehanz):

Or decreasing gradually, of course...

OpenStudy (amistre64):

\[\frac{ae^{100}-ae^0}{100-0}=.001\] :) solve for a

OpenStudy (zehanz):

So you have to know something about the b in \(b^x\)

OpenStudy (zehanz):

@dumbcow 's reply says it all :D

OpenStudy (anonymous):

Thanks soo much, guys! :) I appreciate it. :)

OpenStudy (anonymous):

So I created this...would this work? 15*(1.4)^x

OpenStudy (zehanz):

1.4 >1, so yes, this will do the job!

OpenStudy (anonymous):

Thanks soo much! :)

OpenStudy (anonymous):

@ZeHanz I appreciate it. :)

OpenStudy (zehanz):

YW!